# Pushing the frontiers of subexponential FPT time for Feedback Vertex Set

### v1

@@ 5. Applications
In this section, we prove that $s$-string graphs and pseudo-disk graphs are nice graph classes for some parameters. We recall that a system of pseudo-disks is a collection of regions in the plane homeomorphic to a disk such that any two of them share at most 2 points of their boundary. Similar arguments are used for both considered classes, but in the case of pseudo-disks the arguments are a bit simpler, we then consider this class in first.
@@
A very important aspect in this result is that it is not required that $\mathcal{F}\cup\mathcal{B}$ is a family of pseudo-disks, and thus pseudo-disks of $\mathcal{B}$ may “cross” pseudo disks of $\mathcal{F}$. This is indeed the case in our applications where in particular we associate to each tree in $G-M$ a pseudo-disk in $\mathcal{B}$, and this pseudo disk may cross pseudo-disks associated to vertices of $\mathcal{M}$ (see proof of 5.4).
@@ 5.1. Application to pseudo-disk graphs
In this section, we prove that the class of pseudo-disk graphs is nice, and so by our main theorem it admits a robust subexponential FPT algorithm for FVS. Note that for this graph class, the existence of such algorithm was revy recently given in [BBGR25].
@@
Let $r\geq 2$, $G$ be a $K_{r,r}$-free pseudo-disk graph, $(\mathcal{D}_{v})_{v\in V(G)}$ be a representation of $G$ as pseudo-disks, $A\subseteq V(G)$, and $\mathcal{T}$ a family of disjoint non-adjacent trees of $G-A$. We would like to apply Theorem 5.1 on the set $A$ and the family of subsets of the plane obtained by taking for each tree $T\in\mathcal{T}$ the union $\mathcal{D}_{T}=\cup_{v\in T}\mathcal{D}_{v}$. Because we are considering pseudo-disks and trees, this union is homeomorphic to a disk. Moreover as the trees are non-adjacent the obtained unions do not intersect each other, and so are a pseudo-disk system trivially. Now applying Theorem 5.1 with the families $(\mathcal{D}_{v})_{v\in A}$ and $(\mathcal{D}_{T})_{T\in\mathcal{T}}$ directly gives that there exists a constant $c_{2}$ such that if for all $T\in\mathcal{T}$ we have $d_{A}(T)\leq p$ then |$|\{N_{A}(T),~{}T\in\mathcal{T}\}|\leq c_{2}p^{3}|A|$ as wanted for the second bound. Observing that we can take $p=|A|$ as an upper bound for $d_{A}(T)$ with $T\in\mathcal{T}$ gives the first wanted bound. ∎
@@
There exist constants $c_{1},c_{2},c_{3},c_{4}$ such that the class of pseudo-disk graphs is a nice class with parameters $\alpha=4,~{}f_{1}(r)=c_{1},~{}f_{2}(r,p,m)=c_{2}p^{3}$, $\delta=\frac{1}{2}$, $f_{r}=c_{3}\sqrt{r\log r}$ and $d_{r}=c_{4}r\log r$.
@@
Most of the conditions were already stated with Theorem 1.2, Theorem 1.3 and 5.2. The only property that remains to check is that in a pseudo-disk graph $G$, contracting an edge between degree-two vertices $u$ and $v$ that do not belong to a triangle, obtaining a new vertex $w$, preserves being a pseudo-disk graph. By considering $(\mathcal{D}_{q})_{q\in V(G)}$ a pseudo-disk representation of $G$, we can take $\mathcal{D}_{w}=\mathcal{D}_{u}\cup\mathcal{D}_{v}$ as it is homeomorphic to a disk and its boundary will cross the second neighbor of $u$ (respectively $v$) as many time than the boundary of $\mathcal{D}_{u}$ (respectively $\mathcal{D}_{v}$). ∎
And so applying our main theorem we obtain: See 1.7 Note that our main theorem gives that it suffices to take $\eta>\frac{44}{45}$ for the result to hold. Sharper results already exist in the literature for pseudo-disk graphs, by generalizing the robust algorithm of [LPS+22] dealing with disk graphs to pseudo-disk graphs, or by using the representation as pseudo-disks for obtaining an even better running time. See [BBGR24b], the full version of [BBGR25], for both methods.
@@ 5.2. Application to $s$-string graphs
The proof is very similar to the proof of 5.2 with some additional tweaks. Let $G$ be a $K_{r,r}$-free $s$-string graph for some $r\geq 2$ and $(\mathcal{S}_{v})_{v\in V(G)}$ be an $s$-string representation of $G$. For $T\in\mathcal{T}$, again we consider the region $\mathcal{D}_{T}=\bigcup_{v\in V(T)}\mathcal{S}_{v}$. In this case however, even so $T$ is a tree, the complement of this region is not necessarily connected (for example if two strings of $T$ intersect on several points). In order to obtain a region of the plane homeomorphic to a disk, we modify $\mathcal{D}_{T}$ by cutting small sections where no intersections take place (not even with a string of $A$) and obtain a new set $\mathcal{D}_{T}^{\prime}\subseteq\mathcal{D}_{T}$ that is connected and such that there is exactly one arc-connected region in $\mathbb{R}^{2}\setminus\mathcal{D}_{T}^{\prime}$ and which intersects the same strings in $A$ as $\mathcal{D}_{T}$ (see Figure 5). Moreover we add a small thickness to $\mathcal{D}_{T}^{\prime}$ while preserving the wanted property about the number of regions in $\mathbb{R}^{2}\setminus\mathcal{D}_{T}^{\prime}$ in order to obtain a geometric shape homeomorphic to a disk. Observe that as there are no edges between vertices of distinct members of $\mathcal{T}$, by taking the thickness small enough the connected sets of the form $\mathcal{D}_{T}^{\prime}$ (for $T\in\mathcal{T}$) are disjoint and so they trivially form a system of pseudo-disks.
@@
Now we would like to construct another system of pseudo-disks representing the vertices of $A$. We do this by cutting the strings in $(\mathcal{S}_{u})_{u\in A}$ whenever two of them cross each other. Observe that by Theorem 1.3 there exists a constant $c$ such that by denoting $d_{r}=cr\log r$ we have at most $d_{r}|A|$ intersecting pairs, and remember that each pair intersects each other at most $s$ times as we consider a $s$-string representation. So we have at most $sd_{r}|A|$ cuts which gives an upper bound of $(sd_{r}+1)|A|$ for the number of sections of strings between intersection points. We denote $C$ the set of such sections. Again we define a family $(\mathcal{R}_{v})_{v\in C}$ by considering the sections (which are disjoint by construction) and giving them some thickness, while preventing the sections to intersect each other. This is clearly a system of pseudo-disks as its elements are not crossing. By Theorem 5.1, the number of distinct sets $N_{C}(\mathcal{D}^{\prime}_{T})=\{v\in C,~{}{\mathcal{D}^{\prime}_{T}}\text{ intersects }{\mathcal{R}_{v}}\}$, for $T\in\mathcal{T}$, is $\mathcal{O}(x^{3}|C|)$ where $x$ is the maximum, over every $T\in\mathcal{T}$, of the number of pseudo-disks in $(\mathcal{R}_{v})_{v\in C}$ crossed by $D^{\prime}_{T}$. This upper bound gives the second result of this lemma by observing that if a pair of pseudo-disks $\mathcal{D}^{\prime}_{T_{1}},\mathcal{D}^{\prime}_{T_{2}}$ intersect exactly the same set of ${\mathcal{R}}_{v}$, then $N_{A}(T_{1})=N_{A}(T_{2})$. We obtain the first result of the lemma by using the trivial bound $x\leq|C|$. ∎
@@
There exist constants $c_{1},c_{2},c_{3},c_{4}$ such that for every $s\geq 1$ the class of $s$-string graphs is a nice class with parameters $\alpha=4,~{}f_{1}(r)=c_{1}s^{4}r\log r,~{}f_{2}(r,p,m)=c_{2}(sr\log r)^{4}(p+m)^{3}$, $\delta=\frac{1}{2}$, $f_{r}=c_{3}\sqrt{r\log r}$ and $d_{r}=c_{4}r\log r$.
@@
Again most of the conditions were already stated with Theorem 1.2, Theorem 1.3 and 5.4. The only remaining property to prove is that given an $s$-string graph $G$, contracting an edge between degree-two vertices $u$ and $v$ that do not belong to a triangle preserves being a $s$-string graph. We do this by replacing in a $s$-string representation $(\mathcal{S}_{q})_{q\in V(G)}$ of $G$ the strings $\mathcal{S}_{u}$ and $\mathcal{S}_{v}$ with a new string $\mathcal{S}_{w}$ while preserving the global property of being an $s$-string family. $\mathcal{S}_{u}$ contains a Jordan arc $a_{u}$ linking its two neighbors, without being intersected by any string in its interior. Let denote $p_{v}$ the extremity of $a_{u}$ in $\mathcal{S}_{v}$. Then $\mathcal{S}_{v}$ contains a Jordan arc $a_{v}$ linking $p_{v}$ to the string representing its second neighbor, without being intersected by any other string than $\mathcal{S}_{u}$. The string $\mathcal{S}_{w}$ is then obtained by taking the union of $a_{u}$ and $a_{v}$. This construction ensures that the number of crossings between two strings other than $\mathcal{S}_{u}$ and $\mathcal{S}_{v}$ of the initial representation remain the same. We now have to bound the number of crossing of $\mathcal{S}_{w}$ with others strings. Recalling that $u$ and $v$ do not share a common neighbor, and that by construction $w$ has $2$ crossings (one for each distinct neighbor), $w$ does not have more than $1$ crossings with another string. So after replacing $\mathcal{S}_{u}$ and $\mathcal{S}_{v}$ by $\mathcal{S}_{w}$ we still have a $s$-string family. ∎
@@ 6. Conclusion
In this paper we gave general sufficient conditions for the existence of subexponential parameterized algorithms for the Feedback Vertex Set problem. Our main theorem unifies the previously known results on several classes of graphs such as planar graphs, map graphs, unit-disk graphs, disk graphs, or more generally pseudo-disk graphs, and string graphs of bounded edge-degree. It also applies to classes where such algorithms were not known to exist, notably intersection graphs of thin objects such as segment graphs and more generally $s$-string graphs.
@@
  • Intersection graphs of objects in higher dimensions. However, as proved in [FLS18], intersection graphs of unit balls in $\mathbb{R}^{3}$ do not admit a subexponential parameterized algorithm for FVS, unless $P=NP$.
  • Graph classes excluding an induced minor. [^7] A graph $H$ is said to be an induced minor of a graph $G$ if it can be obtained from $G$ by deleting vertices and contracting edges. Otherwise, $G$ is said to exclude $H$ as induced minor or to be $H$-induced minor free. This is more general than string graphs (which exclude a subdivided $K_{5}$ as an induced minor). In such classes, Korhonen and Lokshtanov [KL24] recently provided an algorithm solving FVS in time $2^{\mathcal{O}_{H}(n^{2/3}\log n)}$, that is, subexponential in the input size $n$.
Less general than the previous item is the class of string graphs, which is nevertheless the most general class of intersection graphs of (arc-connected) objects in the Euclidean plane. At the time of the writing, it is not excluded that it could be a nice class. So far, we are still missing the property about the bounded tree neighborhood complexity. Note that in our proof for $s$-strings, the bound on the number $s$ of crossing is only used to prove that the class satisfies this property of bounded tree neighborhood complexity. Obtaining a similar result without using the bound on the number of crossings would thus yield the desired generalization. But an even simpler way could be to prove that the general case of string graphs can be reduced to the case of $s$-strings graphs for some “small” $s$ function of the problem parameter $k$. A first idea would be to prove that $K_{t,t}$-free string graphs are $t^{c}$-string graphs for some constant $c$, however this fails as there are string graphs with $n$ vertices, and so which are $K_{n,n}$-free, that require $2^{cn}$ crossings for some constant $c$ [KM91].

### v2

@@ 5. Applications
In this section, we prove that $s$-string graphs and pseudo-disk graphs are nice graph classes for some parameters. We recall that a system of pseudo-disks is a collection of regions in the plane homeomorphic to a disk such that any two of them share at most 2 points of their boundary. Similar arguments are used for both considered classes, but in the case of pseudo-disks the arguments are a bit simpler, we then consider this class first.
@@
A very important aspect in this result is that it is not required that $\mathcal{F}\cup\mathcal{B}$ is a family of pseudo-disks, and thus pseudo-disks of $\mathcal{B}$ may “cross” pseudo disks of $\mathcal{F}$. This is indeed the case in our applications where in particular we associate to each tree in GMG-M a pseudo-disk in $\mathcal{B}$, and this pseudo disk may cross pseudo-disks associated to vertices of $\mathcal{M}$ (see proof of 5.4).
@@ 5.1. Application to pseudo-disk graphs
In this section, we prove that the class of pseudo-disk graphs is nice, and so by our main theorem it admits a robust subexponential FPT algorithm for FVS. Note that for this graph class, the existence of such algorithm was very recently given in [FVS-WG].
@@
Let $r\geq 2$, $G$ be a $K_{r,r}$-free pseudo-disk graph, $(\mathcal{D}_{v})_{v\in V(G)}$ be a representation of $G$ as pseudo-disks, $A\subseteq V(G)$, and $\mathcal{T}$ a family of disjoint non-adjacent trees of GAG-A. We would like to apply Theorem 5.1 on the set $A$ and the family of subsets of the plane obtained by taking for each tree $T\in\mathcal{T}$ the union $\mathcal{D}_{T}=\cup_{v\in T}\mathcal{D}_{v}$. Because we are considering pseudo-disks and trees, this union is homeomorphic to a disk. Moreover as the trees are non-adjacent the obtained unions do not intersect each other, $(\mathcal{D}_{T})_{T\in\mathcal{T}}$ is trivially a system of pseudo-disks. Now applying Theorem 5.1 with the families $(\mathcal{D}_{v})_{v\in A}$ and $(\mathcal{D}_{T})_{T\in\mathcal{T}}$ directly gives that there exists a constant $c_{2}$ such that if for all $T\in\mathcal{T}$ we have $d_{A}(T)\leq p$ then $|\{N_{A}(T),~T\in\mathcal{T}\}|\leq c_{2}p^{3}|A|$ as wanted for the second bound. Observing that we can take $p=|A|$ as an upper bound for $d_{A}(T)$ with $T\in\mathcal{T}$ gives the first wanted bound. ∎
@@
There exist constants $c_{1},c_{2},c_{3},c_{4}$ such that the class of pseudo-disk graphs is a nice class with parameters $\alpha=4,~f_{1}(r)=c_{1},~f_{2}(r,p,m)=c_{2}p^{3}$, $\delta=\frac{1}{2}$, $\tau(r)=c_{3}\sqrt{r\log r}$ and $d(r)=c_{4}r\log r$.
@@
Most of the conditions were already stated with Theorem 1.3, Theorem 1.2 and 5.2. The only property that remains to check is that in a pseudo-disk graph $G$, the contraction of an edge between two degree-two vertices $u$ and $v$ that do not belong to a triangle results in a pseudo-disk graph. By considering $(\mathcal{D}_{q})_{q\in V(G)}$ a pseudo-disk representation of $G$, we can take $\mathcal{D}_{w}=\mathcal{D}_{u}\cup\mathcal{D}_{v}$ as it is homeomorphic to a disk and its boundary will cross the neighbor of $u$ (respectively $v$) as many times as does the boundary of $\mathcal{D}_{u}$ (respectively $\mathcal{D}_{v}$). ∎
And so applying our main theorem we obtain: See 1.7 Note that our main theorem gives that it suffices to take $\eta>\frac{44}{45}$ for the result to hold. Sharper results already exist in the literature for pseudo-disk graphs, by generalizing the robust algorithm of [lokSODA22] dealing with disk graphs to pseudo-disk graphs, or by using the representation as pseudo-disks for obtaining an even better running time. See [preprintFVSpseudo], the full version of [FVS-WG], for both methods.
@@ 5.2. Application to $s$-string graphs
The proof is very similar to the proof of 5.2 with some additional tweaks. Let $G$ be a $K_{r,r}$-free $s$-string graph for some $r\geq 2$ and $(\mathcal{S}_{v})_{v\in V(G)}$ be an $s$-string representation of $G$. For $T\in\mathcal{T}$, again we consider the region $\mathcal{D}_{T}=\bigcup_{v\in V(T)}\mathcal{S}_{v}$. In this case however, even though $T$ is a tree, the complement of this region is not necessarily connected (for example if two strings of $T$ intersect on several points). In order to obtain a region of the plane homeomorphic to a disk, we modify $\mathcal{D}_{T}$ by cutting small sections where no intersections take place (not even with a string of $A$) and obtain a new set $\mathcal{D}_{T}^{\prime}\subseteq\mathcal{D}_{T}$ that is connected and such that there is exactly one arc-connected region in $\mathbb{R}^{2}\setminus\mathcal{D}_{T}^{\prime}$ and which intersects the same strings in $A$ as $\mathcal{D}_{T}$ (see Figure 5). Moreover we add a small thickness to $\mathcal{D}_{T}^{\prime}$ while preserving the wanted property about the number of regions in $\mathbb{R}^{2}\setminus\mathcal{D}_{T}^{\prime}$ in order to obtain a geometric shape homeomorphic to a disk. Observe that as there are no edges between vertices of distinct members of $\mathcal{T}$, by taking the thickness small enough the connected sets of the form $\mathcal{D}_{T}^{\prime}$ (for $T\in\mathcal{T}$) are disjoint and so they trivially form a system of pseudo-disks.
@@
Now we would like to construct another system of pseudo-disks representing the vertices of $A$. We do this by cutting the strings in $(\mathcal{S}_{u})_{u\in A}$ whenever two of them cross each other. Observe that by Theorem 1.2 there exists a constant $c$ such that by denoting $d(r)=cr\log r$ we have at most $d(r)|A|$ intersecting pairs, and remember that each pair intersects each other at most $s$ times as we consider a $s$-string representation. So we have at most $sd(r)|A|$ cuts which gives an upper bound of $(sd(r)+1)|A|$ for the number of sections of strings between intersection points. We denote $C$ the set of such sections. Again we define a family $(\mathcal{R}_{v})_{v\in C}$ by considering the sections (which are disjoint by construction) and giving them some thickness, while preventing the sections to intersect each other. This is clearly a system of pseudo-disks as its elements are not crossing. By Theorem 5.1, the number of distinct sets $N_{C}(\mathcal{D}^{\prime}_{T})=\{v\in C,~{\mathcal{D}^{\prime}_{T}}\text{ intersects }{\mathcal{R}_{v}}\}$, for $T\in\mathcal{T}$, is $\mathcal{O}(x^{3}|C|)$ where $x$ is the maximum, over every $T\in\mathcal{T}$, of the number of pseudo-disks in $(\mathcal{R}_{v})_{v\in C}$ crossed by $D^{\prime}_{T}$. This upper bound gives the second result of this lemma by observing that if a pair of pseudo-disks $\mathcal{D}^{\prime}_{T_{1}},\mathcal{D}^{\prime}_{T_{2}}$ intersect exactly the same set of ${\mathcal{R}}_{v}$, then $N_{A}(T_{1})=N_{A}(T_{2})$. We obtain the first result of the lemma by using the trivial bound $x\leq|C|$. ∎
@@
There exist constants $c_{1},c_{2},c_{3},c_{4}$ such that for every $s\geq 1$ the class of $s$-string graphs is a nice class with parameters $\alpha=4,~f_{1}(r)=c_{1}s^{4}r\log r,~f_{2}(r,p,m)=c_{2}(sr\log r)^{4}(p+m)^{3}$, $\delta=\frac{1}{2}$, $\tau(r)=c_{3}\sqrt{r\log r}$ and $d(r)=c_{4}r\log r$.
@@
Again most of the conditions were already stated with Theorem 1.3, Theorem 1.2 and 5.4. The only remaining property to prove is that given an $s$-string graph $G$, contracting an edge between degree-two vertices $u$ and $v$ that do not belong to a triangle preserves being a $s$-string graph. We do this by replacing in a $s$-string representation $(\mathcal{S}_{q})_{q\in V(G)}$ of $G$ the strings $\mathcal{S}_{u}$ and $\mathcal{S}_{v}$ with a new string $\mathcal{S}_{w}$ while preserving the global property of being an $s$-string family. $\mathcal{S}_{u}$ contains a Jordan arc $a_{u}$ linking its two neighbors, without being intersected by any string in its interior. Let denote $p_{v}$ the extremity of $a_{u}$ in $\mathcal{S}_{v}$. Then $\mathcal{S}_{v}$ contains a Jordan arc $a_{v}$ linking $p_{v}$ to the string representing its second neighbor, without being intersected by any other string than $\mathcal{S}_{u}$. The string $\mathcal{S}_{w}$ is then obtained by taking the union of $a_{u}$ and $a_{v}$. This construction ensures that the number of crossings between two strings other than $\mathcal{S}_{u}$ and $\mathcal{S}_{v}$ of the initial representation remain the same. We now have to bound the number of crossing of $\mathcal{S}_{w}$ with others strings. Recalling that $u$ and $v$ do not share a common neighbor, and that by construction $w$ has $2$ crossings (one for each distinct neighbor), $w$ does not have more than one crossing with some other string. So after replacing $\mathcal{S}_{u}$ and $\mathcal{S}_{v}$ by $\mathcal{S}_{w}$ we still have a $s$-string family. ∎
@@ 6. Conclusion
In this paper we gave general sufficient conditions for the existence of subexponential parameterized algorithms for the Feedback Vertex Set problem. Our main theorem unifies the previously known results on several classes of graphs such as planar graphs, map graphs, unit-disk graphs, disk graphs, or more generally pseudo-disk graphs, and string graphs of bounded degree. It also applies to classes where such algorithms were not known to exist, notably intersection graphs of thin objects such as segment graphs and more generally $s$-string graphs.
@@
  • Intersection graphs of objects in higher dimensions. However, as proved in [fomin2018excluded], under ETH intersection graphs of unit balls in $\mathbb{R}^{3}$ do not admit a subexponential parameterized algorithm for FVS.
  • Graph classes excluding an induced minor. [^9] A graph $H$ is said to be an induced minor of a graph $G$ if it can be obtained from $G$ by deleting vertices and contracting edges. Otherwise, $G$ is said to exclude $H$ as induced minor or to be $H$-induced minor free. This is more general than string graphs (which exclude a subdivided $K_{5}$ as an induced minor). In such classes, Korhonen and Lokshtanov [korhonen2023induced] recently provided an algorithm solving FVS in time $2^{\mathcal{O}_{H}(n^{2/3}\log n)}$, that is, subexponential in the input size $n$. So far a subexponential FPT algorithm (parameterized by $k$) is still missing.
Less general than the previous item is the class of string graphs, which is nevertheless the most general class of intersection graphs of (arc-connected) objects in the Euclidean plane. This class is not a nice class, as we are missing the property of bounded tree neighborhood complexity. In fact, even the more restricted case of the class of outerstring graphs [^10] Outerstring graphs are the string graphs with a representation such that there is a circle containing at least one endpoint of each string. does not have bounded (tree) neighborhood complexity. This is witnessed by a construction we describe now, that is illustrated in Figure 7. For $r>3$, construct an outerstring graph $G$ as follows: consider a clique of size $r$ obtained as an outerstring graph by taking $r$ strings with one extremity on a circle and then permuting the order of the strings in order to obtain the $r!$ permutations. It is then not hard to see that it is possible to add $2^{r}$ disjoint strings with one endpoint on the circle all crossing distinct subsets of the $r$ initial strings. The graph $G$ is then the intersection graph of those $r+2^{r}$ strings. Observe that $G$ has only $r+1$ vertices with degree at least $r$, and so does not have $K_{r,r}$ as a subgraph. By taking $A\subseteq V(G)$ the vertices associated to the $r$ first strings, we have $|\{N(v)\cap A:v\in V(G)\}|=2^{r}$. A bound of the form $\mathcal{O}(r^{c}\cdot|A|)$ for some $c$ is thus impossible. Note that we can create examples of arbitrarily large size by taking disjoint copies of this construction.
Figure 7. The class of outerstring graphs does not have bounded tree neighborhood complexity, as in a set $A$ of $t$ outerstrings crossing each other in a way their top to bottom order goes through many permutations (the strings in color in the figure, with $t=4$), it is possible to generate $2^{t}$ distinct neighborhoods in AA (in black).
As 3.2 allows reducing to $K_{t,t}$-free subgraphs, one way to deal with string graphs would be proving that $K_{t,t}$-free string graphs are $t^{c}$-string graphs for some constant $c$. Unfortunately, this fails as there are string graphs with $n$ vertices, hence $K_{n,n}$-free, that require $2^{cn}$ crossings for some constant $c$ [stringexpo].